A number, half of that number, and one-third of that number are added. The result is 22. What is the original number?
Answer by checkley75(3666)

X+X/2+X/3=22
(6X+3X+2X)/6=22
11X/6=22 NOW CROSS MULTIPLY
11X=132
X=132/11
X=12
PROOF
12+12/2+12/3=22
12+6+4=22
22=22